OpenStudy (aravindg):

Find domain and range of function f(x)=[(x^{2}-3x+2/x^{2}+x-6]

OpenStudy (anonymous):

Okay Maybe \[ f(x) = \frac{x^2 -3x +2}{x^2+x-6}\] Right ?

OpenStudy (aravindg):

yes

OpenStudy (anonymous):

Now I will advise Factorization

OpenStudy (aravindg):

plz do it and giv answer i am confused

OpenStudy (anonymous):

Okay \[f(x) =\frac{(x-1)(x-2)}{(x+3)(x-2)}\] Now ...\[x=\neq -3,2\]

OpenStudy (anonymous):

So that makes Domain \[x \in{R} - (-3,2)\]

OpenStudy (anonymous):

Now to Find Range ..... \[Let\:\:\:\:\:\:\:\f(x) = y\] \[y = \frac{x-1}{x+3}\]

OpenStudy (aravindg):

thx a lot plzz answer my othr 2 questions

OpenStudy (anonymous):

\[x = \frac{1+y}{1-y}\] That makes range \[y \in R - (1)\]

OpenStudy (aravindg):

thx a lot plzz answer my othr 2 questions o

OpenStudy (anonymous):

the domain is okay, the range is okay too, except the experession of x in terms of y, I think as follows: x=\[x=(3y+1)/(1-y)\]